OpenStudy (anonymous):

I'd be very grateful for help! Solve the following system of equations. Please show your work to receive full credit. x - y = 19 2x + y = 2

jimthompson5910 (jim_thompson5910):

if you add the equations, you get this x - y = 19 2x + y = 2 ---------- 3x + 0y = 21 3x = 21 x = ??

OpenStudy (anonymous):

oh wow 7?

jimthompson5910 (jim_thompson5910):

so if x = 7, then what is y

OpenStudy (anonymous):

so I plug in 7 for one of the equations?

jimthompson5910 (jim_thompson5910):

correct, pick any equation, replace x with 7, then solve for y

OpenStudy (anonymous):

unsure if this is right, but 12? if it isnt I know why.

jimthompson5910 (jim_thompson5910):

let's pick the first equation

jimthompson5910 (jim_thompson5910):

x - y = 19 7 - y = 19 ... plug in x = 7 -y = 19 - 7 -y = 12 -1y = 12 y = 12/(-1) y = -12

jimthompson5910 (jim_thompson5910):

so you were off by a sign

OpenStudy (anonymous):

Oh! I rushed it a tad, that's a simple fix! I appreciate the help you are a god haha :D

OpenStudy (anonymous):

and is this substitution method?

jimthompson5910 (jim_thompson5910):

no it's the elimination method

jimthompson5910 (jim_thompson5910):

go back up the very top and you'll see that we're adding the equations and one variable is eliminated that way

OpenStudy (anonymous):

oh wow I should of understood that...

jimthompson5910 (jim_thompson5910):

the substitution method would have us isolate a variable, then substitute or plug it into the other equation

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

just keep practicing and you'll get a better hang of things

OpenStudy (anonymous):

Im doing a few practice problems, Using the method you have shown me:)

jimthompson5910 (jim_thompson5910):

I'm glad it's clicking

OpenStudy (rina.r):